there are n students in a class a group of 3 goes to appughar with teacher and the same set of children doesn't go more than once .it is found that the teacher goes to appughar  84times more than a particular child find n?????


ans  n=10


 

Grade:11

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Akhilesh

no. of times teacher goes to appughar = nC3

no. of times a particular child goes to appughar = (n-1)C2      bcoz we have to select 2 students out of n-1 students

According to question

nc3-(n-1)C2=84

after solving

(n-1)(n-2)(n-3)=504

n=10

 

All the best

AKASH GOYAL
AskiitiansExpert-IITD

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.